The sum of two consective whole numbers is 57. What are the two numbers? (Hint: Let n represent the first number. Then n +1 is the next consecutive whole number.)
Answer by mananth(16946)

n + n+1 =57
2n+1=57
-1
2n+1-1=57-1
2n=56
/2
n=56/2
n=28
..
the numbers are 28,29